# Evolution of a Long-Lived Deep-Seated Main-Sequence Magnetic Field During White Dwarf Cooling

### v1

@@ 1 Introduction
Magnetic fields have been consistently detected on the surface of WDs for the past 55 years (Kemp et al., 1970; Angel & Landstreet, 1970), and today more than 600 magnetized WDs are known. These fields have been observed across a diverse range of WD types, including both single and binary systems, spanning various spectral classes, masses, and evolutionary stages (see Ferrario et al., 2015, 2020, for reviews). The strength of these magnetic fields varies widely, from $10^{3}$ to $10^{9}$ G. Recent studies of the volume-limited sample of magnetic WDs within 20 parsecs of the Sun have contributed significantly to our understanding of these objects (Bagnulo & Landstreet, 2021; Kawka et al., 2007). These studies have systematically examined each of the WDs in this region searching for magnetic fields, thus eliminating the observational biases inherent in previous magnitude-limited surveys. Their findings reveal that magnetic WDs tend to be, on average, more massive than their non-magnetic counterparts. Moreover, they discovered a strong correlation between magnetism and core crystallization, with a higher prevalence of magnetic fields in WDs that have undergone or are currently undergoing the crystallization process. Further expanding on this work, Bagnulo & Landstreet (2022) extended the sample to 40 parsecs, but only considering those WDs younger than 0.6 gigayears (Gyr). Their analysis reaffirmed previous conclusions, confirming that the occurrence of magnetism in young (pre-crystallization) WDs is approximately 10%, whereas it increases to around 30% in older, crystallized WDs.
@@
A recent alternative explanation for the WD magnetic fields has been explored in Camisassa et al. (2024), that proposes that magnetic fields arise as a result of the surface emergence of a deep-seated convective dynamo action. Main sequence stars more massive than $\sim 1.1M_{\odot}$ have convective cores that can sustain dynamo magnetic fields. This idea is supported by magnetohydrodynamic (MHD) simulations of convective cores of main sequence stars (e.g. Augustson et al., 2016; Hidalgo et al., 2024) and by the astroseismologic detection of magnetic fields buried in the deep interior of red giant stars (e.g. Stello et al., 2016; Deheuvels et al., 2023). In an exploratory paper, Camisassa et al. (2024) studied the evolution of a main-sequence dynamo magnetic field through the subsequent evolutionary phases to the white dwarf phase. These authors found that, if this field is generated and can survive trapped in the stellar interior throughout the evolution, it would emerge to the surface during the WD stage. Moreover, the time of surface emergence of the main-sequence dynamo can explain most of the magnetic WDs with masses $\gtrsim 0.65M_{\odot}$. However, this work relies on a basic estimation of the time of emergence through magnetic diffusion, and this basic picture cannot predict the magnetic field strength at the WD surface.
@@
2 Methods
2.1 Main sequence dynamo and magnetic boundaries
@@
2.2 White dwarf models
@@
Figure 1: Magnetic field strength of the main sequence dynamo assuming equipartition for main sequence stars of 1.5, 3, 4, and 5 M. The solid vertical lines indicate the final white dwarf masses, assuming the initial-to-final-mass relation of Catalán et al. (2008). The dashed vertical lines indicate the location of the magnetic boundaries below which we expect the magnetic field to be buried at the beginning of the white dwarf phase, assuming the cubic fit prescription from Camisassa et al. (2024). Note that in the x-axis $\mathrm{M_{r}}=m/M_{\odot}$, where $m$ is the total mass in solar masses contained at the location measured from the center.
2.3 Magnetic Diffusion Calculation
@@
$\frac{\partial{\mbox{{\boldmath$B$}}}}{\partial{t}}=-{\mbox{{\boldmath$\nabla$}}}\times(\eta{\mbox{{\boldmath$\nabla$}}}\times{\mbox{{\boldmath$B$}}}),$ (1)
where $\eta$ is the magnetic diffusivity. Using the WD models described in Section 2.2 and the Electron Transport Coefficients of Magnetized Stellar Plasmas code of Potekhin et al. (1999, 2015), we computed the electrical conductivities $\sigma$ along the star radial coordinate for different cooling times since the formation of the WD. Thus, in equation 1 we include the Ohmic diffusivity as $\eta_{\mathrm{ohm}}=c^{2}/4\pi\sigma$, where $c$ is the speed of light. As presented in Castro-Tapia et al. (2024b), we used a free parameter $\mathrm{f_{Rm}}$ proportional to the magnetic Reynolds number to calculate the effect of turbulent diffusion as $\eta_{\mathrm{turb}}=\mathrm{f_{Rm}}\eta_{\mathrm{ohm}}$, and then the full conductivity including both contributions is $\eta=\eta_{\mathrm{ohm}}+\eta_{\mathrm{turb}}$.
Assuming an axisymmetric poloidal field and taking a separable magnetic vector potential ${\mbox{{\boldmath$A$}}}=A_{\phi}(r,\theta,t)\hat{e}_{\phi}=\sum_{l}[R_{l}(r,t)/r]P^{1}_{l}(\cos{\theta})\hat{e}_{\phi}$, where $R_{l}$ is the radial component and $P^{1}_{l}$ is the associated Legendre polynomial of order 1 (as done in Castro-Tapia et al. 2024b, motivated by previous work of Cumming 2002 and references therein), we use ${\mbox{{\boldmath$B$}}}={\mbox{{\boldmath$\nabla$}}}\times{\mbox{{\boldmath$A$}}}$ to write
@@
Thus, to obtain the diffusion of the magnetic field for our WD models with surviving fields from the core-convective dynamo, we solve numerically equation 2 using the Crank–Nicolson method taking as initial condition that the field is filling the WD interior up to the magnetic boundary computed from the fit of Camisassa et al. (2024).
3 Magnetic field evolution
3.1 Dipole and quadrupole components
@@
${\mbox{{\boldmath$B_{\mathrm{d}}$}}}(r,\theta,t)=\frac{2R_{1}(r,t)}{r^{2}}\cos{\theta}\hat{e}_{r}+\frac{1}{r}\frac{\partial R_{1}(r,t)}{\partial{r}}\sin{\theta}\hat{e}_{\theta},$ (3)
@@
${\mbox{{\boldmath$B_{\mathrm{d+q}}$}}}(r,\theta,t)=\frac{1}{r^{2}}\left[2R_{1}(r,t)\cos{\theta}+3R_{2}(r,t)(3\cos^{2}{\theta}-1)\right]\hat{e}_{r}$
@@
We use the boundary conditions $\partial R_{l}/\partial r=(l+1)R_{l}/r$ for the center of the star $r\rightarrow{0}$ and $\partial R_{l}/\partial r=-lR_{l}/r$ for $r\rightarrow{R_{\mathrm{WD}}}$, which enforces that the current fulfill ${\mbox{{\boldmath$J$}}}\propto{\mbox{{\boldmath$\nabla$}}}\times{\mbox{{\boldmath$B$}}}=0$ outside the star (Cumming, 2002).
@@
$B_{\mathrm{rms}}(r,t_{0})=\sqrt{\frac{1}{4\pi}\int_{\Omega}|{\mbox{{\boldmath$B$}}}|^{2}(r,\theta,t_{0})d\Omega}$
@@
and for ${\mbox{{\boldmath$B_{\mathrm{d+q}}$}}}(r,\theta,t_{0})$ we also considered the initial condition $R_{1}(r,t_{0})=R_{2}(r,t_{0})$ for simplicity.
@@
Figure 3: Evolution of the magnetic field strength (top panel) and the magnetic diffusivity (bottom) as a function of the normalized white dwarf radius for our 0.6 M model, assuming an initial axisymmetric dipole field.
@@
3.2 Magnetic field evolution for different masses
We computed the magnetic field evolution using the methodology described in Section 2.3 for the different WD models evolved with MESA. For each mass, we considered four cases of the magnetic field evolution to cover most of the physical processes expected during WD cooling and the expectations of the magnetic field geometry from observations. Thus, the cases covered two geometries of the field: dipole and dipole+quadrupole, each combined with two cases of diffusivity, only ohmic diffusion, and the other ohmic+turbulent diffusion. The latter case of diffusivity is used to consider turbulent diffusion driven by compositional convection during crystallization (Castro-Tapia et al., 2024b) since many of the observed magnetic WDs with field strengths $\gtrsim 10^{6}$ G have cooling ages larger that the time expected for the onset of crystallization (Bagnulo & Landstreet, 2021, 2022).
In Figure 3 we show a example temporal evolution of the $B_{\mathrm{rms}}$ and diffusivity $\eta$ as a function of the radius for the $0.6\ M_{\odot}$ model assuming a dipole geometry and taking $\mathrm{f_{RM}}=10$ for turbulent diffusivity. This figure is similar to Figure 3 of Castro-Tapia et al. (2024b). However, in the current case, we note how the magnetic field has already diffused to the surface of the WD when less than $5\%$ of the crystal core has formed as a direct consequence of the initial condition of the magnetic boundary computed by Camisassa et al. (2024) and letting the magnetic field evolve from the formation of the WD.
In order to later compare with observations, we computed the surface value of $B_{\mathrm{rms}}$ as a function of time from the magnetic field evolution of the models with different masses. In the left panel of Figure 4, we show $B_{\mathrm{rms,surf}}$ normalized to the initial value $B_{0}$, as a function of the cooling time when considering only ohmic diffusion for the different masses and d and d+q field geometries. We define the reduction factor $f_{B}(t_{\mathrm{cool}})=(B_{\mathrm{rms,sur}}/B_{0})(t_{\mathrm{cool}})$, which indicates the expected value of the magnetic field at the WD surface, normalized by the initial field, as a function of the WD cooling time. In the right panel we show the same but including turbulent diffusion with $\mathrm{f_{RM}}=10$. Regardless of the mass, all the cases show that $B_{\mathrm{rms,surf}}$ increases as it is transported from the interior to the surface, reaching a maximum intensity, to then decay, as it is diffused out of the WD, with a much smaller rate of change than the first increase. This slow decay is a direct consequence of the crystallization process that increases the conductivities as the solid phase grows.
@@ 3.2 Magnetic field evolution for different masses
4 Comparison with observations
Figure 5: Color map indicating the ratio between the surface magnetic field and the initial magnetic field buried below the magnetic boundary, as a function of white dwarf mass and cooling time. The local 20-40pc sample of white dwarfs from Bagnulo & Landstreet (2022) is plotted using orange crosses for non-magnetic white dwarfs and dots for magnetic white dwarfs, where the surrounding circle is proportional to the logarithm of the magnetic field strength.
@@ 4 Comparison with observations
Figure 6: Left panel: Magnetic fields at surface of white dwarf stars taken from the Montreal White Dwarf Database (red crosses), and their corresponding magnetic fields expected at the edge of the magnetic boundary during the main sequence stage, assuming different considerations for the field geometry and diffusivities (circles). The solid lines indicate the magnetic fields expected at this magnetic boundary during the main sequence, taken from the MHD simulations of main sequence A stars of Brun et al. (2005) (green lines) and Hidalgo et al. (2024) (black line) and from our estimations assuming equipartition between the kinetic and magnetic energy (purple line). The blue shaded area indicates the typical magnetic field strengths measured at the surface of magnetic main sequence Ap and Bp stars. Right panel: Same as left panel but for the sample of magnetic white dwarfs within 20pc from our Sun and magnetic white dwarfs younger than 0.6 Gyrs within 40pc from our Sun (Bagnulo & Landstreet, 2022).
@@
Figure 7: Predictions of the surface magnetic field strengths (rms) as a function of cooling age for our models compared to observed magnetic fields in a 100 pc sample of the MWDD. The initial magnetic field prediction $B_{0}$ comes from equipartition in the main-sequence and rescaling assuming conservation of the magnetic flux (Eq. 6). We have separated into different panels the masses of the models in the range $0.6$-$1\ M_{\odot}$, and we compare in each panel with MWDs with masses within a range of $\pm 0.05\ M_{\odot}$ for the corresponding model mass. The solid lines show the d geometry models, while the dashed lines are the d+q models. In the bottom-right panel, we summarize the range of the magnetic field strengths predicted and observed for the different masses; for the observations, the error bars in the vertical axis cover the whole field strength for the given mass bin and have as a data point the logarithmic average of the magnetic fields in the bin, whereas for the model range, the upper limit is the maximum strength predicted, and the arrow is at the minimum $B_{\mathrm{rms,surf}}$ for the age of the youngest WD in the bin.
[…]
@@
In order to predict the value of the magnetic field in the main-sequence phase $B_{\mathrm{MS}}$ necessary to reproduce the observed surface fields in MWDs we consider the reduction factors $f_{B}(t_{\mathrm{cool}})$ from the different evolution models and a scaling of $B_{\mathrm{MS}}=B_{0}(R_{\mathrm{MS,b}}/R_{\mathrm{WD,b}})^{2}$ assuming magnetic flux conservation, where $R_{\mathrm{MS,b}}$ and $R_{\mathrm{WD,b}}$ are the radius at the magnetic boundary of the main-sequence progenitor and the WD, respectively. Thus, for a given observed field $B_{\mathrm{rms,surf}}$ in the WD phase, its corresponding magnetic field strength generated during the main-sequence can be computed as follows
@@
where $f_{B}(t_{\mathrm{cool}})$ is the prediction of our model at a given cooling time $t_{\mathrm{cool}}$ because we do not know $B_{0}$ a priori.
@@
In Figure 6 we show the observed magnetic field in MWDs (red crosses) for different masses and their corresponding predictions of the main-sequence field $B_{\mathrm{MS}}$ (empty circles) necessary to reproduce the observations, obtained using equation 6 for our models with d and d+q geometries. We also included the prediction from our models with turbulent diffusion. In the left panel of the figure, we show the MWDs in the MWDD, while the right panel shows the MWDs in the 20-40 pc sample of Bagnulo & Landstreet (2022). To compare our $B_{\mathrm{MS}}$ predictions, we added the magnetic field strength prediction from a linear fit to the equipartition value at the magnetic boundary of main-sequence progenitors for different WDs, assuming an initial-to-final-mass relation from (Catalán et al., 2008). We also compare the predicted magnetic field strengths from core convection in the main-sequence obtained from MHD simulations of Brun et al. (2005) and Hidalgo et al. (2024), for Main Sequence masses of 2 and 2.2M, respectively.
@@
We note that the necessary $B_{\mathrm{MS}}$ predictions to reproduce the observations are mostly in the range $\sim 10^{3}$-$10^{5}$ G, which is very similar to the equipartition and MHD predictions of $\sim 10^{4}$-$10^{5}$ G. The equipartition prediction appears almost as an upper limit for the main-sequence convective core dynamo field necessary to reproduce the observed MWDs with $\gtrsim 0.75\ M_{\odot}$, except for a very few whose $B_{\mathrm{MS}}$ predictions are larger than the equipartition by a factor of $\lesssim 5$. For MWDs with $\lesssim 0.75\ M_{\odot}$, the $B_{\mathrm{MS}}$ predictions appear more scattered from the equipartition value, some of them even being the highest $B_{\mathrm{MS}}$ for all the masses. This occurs because of the small magnetic boundaries and larger radius for the less massive WDs, which lead to a maximum factor $f_{B}\lesssim 0.03$. Therefore, to reproduce MWDs with about the same few MG strength, the $B_{\mathrm{MS}}$ must be larger for the less massive stars. However, for all the masses, about half of the magnetic fields observed in WDs could have originated from $B_{\mathrm{MS}}$ weaker than $10^{4}$ G, below the equipartition and MHD simulations prediction. Some MWDs with $10^{5}$-$10^{6}$ G could have even been produced with $B_{\mathrm{MS}}$ of just a hundred G. It is important to remark that the initial-to-final-mass relation and, consequently, the value of $\frac{R_{\mathrm{MS,b}}}{R_{\mathrm{WD,b}}}$ are subject to large uncertainties, particularly for the more massive WDs. Therefore, the value of $B_{\mathrm{MS}}$ could be much weaker or stronger than predicted here.

### v2

@@ I Introduction
Magnetic fields have been consistently detected on the surface of WDs for the past 55 years (Kemp et al., 1970; Angel & Landstreet, 1970), and today more than 600 magnetized WDs are known. These fields have been observed across a diverse range of WD types, including both single and binary systems, spanning various spectral classes, masses, and evolutionary stages (see Ferrario et al., 2015, 2020, for reviews). The strength of these magnetic fields varies widely, from $10^{3}$ to $10^{9}$ G. Recent studies of the volume-limited sample of magnetic WDs within 20 parsecs of the Sun have contributed significantly to our understanding of these objects (Kawka et al., 2007; Bagnulo & Landstreet, 2021). These studies have systematically examined each of the WDs in this region searching for magnetic fields, thus eliminating the observational biases inherent in previous magnitude-limited surveys. Their findings reveal that magnetic WDs tend to be, on average, more massive than their non-magnetic counterparts. Moreover, they discovered a strong correlation between magnetism and core crystallization, with a higher prevalence of magnetic fields in WDs that have undergone or are currently undergoing the crystallization process. Further expanding on this work, Bagnulo & Landstreet (2022) extended the sample to 40 parsecs, but only considering those WDs younger than 0.6 gigayears (Gyr). Their analysis reaffirmed previous conclusions, confirming that the occurrence of magnetism in young (pre-crystallization) WDs is approximately 10%, whereas it increases to around 30% in older, crystallized WDs.
@@
A recent alternative explanation for the WD magnetic fields has been explored in Camisassa et al. (2024), which proposes that magnetic fields arise as a result of the surface emergence of a deep-seated convective dynamo action. Main-sequence stars more massive than $\sim 1.1M_{\odot}$ have convective cores that can sustain dynamo magnetic fields. This idea is supported by magnetohydrodynamic (MHD) simulations of convective cores of main sequence stars (e.g. Augustson et al., 2016; Hidalgo et al., 2024) and by the astroseismologic detection of magnetic fields buried in the deep interior of red giant stars (e.g. Stello et al., 2016; Deheuvels et al., 2023). Specifically, Camisassa et al. (2024) studied the evolution of a main-sequence dynamo magnetic field through the subsequent evolutionary phases to the white dwarf phase. These authors found that, if this field is generated and can survive trapped in the stellar interior throughout the evolution, it would emerge to the surface during the WD stage. Moreover, the time of surface emergence of the main-sequence dynamo can explain most of the magnetic WDs with masses $\gtrsim 0.65M_{\odot}$. However, this work relies on a basic estimation of the time of emergence through magnetic diffusion, and this basic picture cannot predict the magnetic field strength at the WD surface.
@@
II Methods
II.1 Main sequence dynamo and magnetic boundaries
@@
II.2 White dwarf models
@@
Figure 1: Magnetic field strength of the main sequence dynamo assuming equipartition for main sequence stars of 1.5, 3, 4, and 5 M. The solid vertical lines indicate the final white dwarf masses, assuming the initial-to-final-mass relation of Catalán et al. (2008). The dashed vertical lines indicate the location of the magnetic boundaries below which we expect the magnetic field to be buried at the beginning of the white dwarf phase, assuming the cubic fit prescription from Camisassa et al. (2024). Note that in the x-axis $\mathrm{M_{r}}=m/M_{\odot}$, where $m$ is the total mass contained at the location from the center in solar masses.
II.3 Magnetic Diffusion Calculation
@@
$\frac{\partial{\mbox{{$B$}}}}{\partial{t}}=-{\mbox{{$\nabla$}}}\times(\eta{\mbox{{$\nabla$}}}\times{\mbox{{$B$}}}),$ (1)
where $\eta$ is the magnetic diffusivity. Using the WD models described in Section II.2 and the Electron Transport Coefficients of Magnetized Stellar Plasmas code of Potekhin et al. (1999, 2015), we computed the electrical conductivities $\sigma$ along the star radial coordinate for different cooling times since the formation of the WD. Thus, in equation (1) we include the Ohmic diffusivity as $\eta_{\mathrm{ohm}}=c^{2}/4\pi\sigma$, where $c$ is the speed of light. As presented in Castro-Tapia et al. (2024b), we used a free parameter $\mathrm{f_{Rm}}$ proportional to the magnetic Reynolds number to calculate the effect of turbulent diffusion as $\eta_{\mathrm{turb}}=\mathrm{f_{Rm}}\eta_{\mathrm{ohm}}$, and then the full conductivity including both contributions is $\eta=\eta_{\mathrm{ohm}}+\eta_{\mathrm{turb}}$.
Assuming an axisymmetric poloidal field and taking a separable magnetic vector potential ${\mbox{{$A$}}}=A_{\phi}(r,\theta,t)\hat{e}_{\phi}=\sum_{l}[R_{l}(r,t)/r]P^{1}_{l}(\cos{\theta})\hat{e}_{\phi}$, where $R_{l}$ is the radial component and $P^{1}_{l}$ is the associated Legendre polynomial of order 1 (as done in Castro-Tapia et al. 2024b, motivated by previous work of Cumming 2002 and references therein), we use ${\mbox{{$B$}}}={\mbox{{$\nabla$}}}\times{\mbox{{$A$}}}$ to write
@@
Thus, to obtain the diffusion of the magnetic field for our WD models with surviving fields from the core-convective dynamo, we solve numerically equation (2) using the Crank–Nicolson method, taking as initial condition that the field is filling the WD interior up to the magnetic boundary computed from the fit of Camisassa et al. (2024).
III Magnetic field evolution
III.1 Dipole and quadrupole components
@@
${\mbox{{$B_{\mathrm{d}}$}}}(r,\theta,t)=\frac{2R_{1}(r,t)}{r^{2}}\cos{\theta}\hat{e}_{r}+\frac{1}{r}\frac{\partial R_{1}(r,t)}{\partial{r}}\sin{\theta}\hat{e}_{\theta},$ (3)
@@
${\mbox{{$B_{\mathrm{d+q}}$}}}(r,\theta,t)=\frac{1}{r^{2}}\left[2R_{1}(r,t)\cos{\theta}+3R_{2}(r,t)(3\cos^{2}{\theta}-1)\right]\hat{e}_{r}$
@@
We use the boundary conditions $\partial R_{l}/\partial r=(l+1)R_{l}/r$ for the center of the star $r\rightarrow{0}$ and $\partial R_{l}/\partial r=-lR_{l}/r$ for $r\rightarrow{R_{\mathrm{WD}}}$, which enforces that the current fulfill ${\mbox{{$J$}}}\propto{\mbox{{$\nabla$}}}\times{\mbox{{$B$}}}=0$ outside the star (Cumming, 2002).
@@
$B_{\mathrm{rms}}(r,t_{0})=\sqrt{\frac{1}{4\pi}\int_{4\pi}|{\mbox{{$B$}}}|^{2}(r,\theta,t_{0})d\Omega}$
@@
and for ${\mbox{{$B_{\mathrm{d+q}}$}}}(r,\theta,t_{0})$ we also considered the initial condition $R_{1}(r,t_{0})=R_{2}(r,t_{0})$ for simplicity.
@@
Figure 3: Evolution of the magnetic field strength (top panel) and the magnetic diffusivity (bottom) as a function of the normalized white dwarf radius for our 0.6 M model, assuming an initial axisymmetric dipole field. In this case, we included turbulent diffusivity inside the compositionally-driven convection zone with $\mathrm{f_{RM}}=10$.
@@
III.2 Magnetic field evolution for different masses
We computed the magnetic field evolution using the methodology described in Section II.3 for the different WD models evolved with MESA. For each mass, we considered four cases of the magnetic field evolution to cover most of the physical processes expected during WD cooling and the expectations of the magnetic field geometry from observations. Thus, the cases covered two geometries of the field: dipole and dipole+quadrupole, each combined with two cases of diffusivity, only ohmic diffusion, and the other ohmic+turbulent diffusion. The latter case of diffusivity is used to consider turbulent diffusion driven by compositional convection during crystallization (Castro-Tapia et al., 2024b) since many of the observed magnetic WDs with field strengths $\gtrsim 10^{6}$ G have cooling ages larger that the time expected for the onset of crystallization (Bagnulo & Landstreet, 2021, 2022).
In Figure 3 we show an example temporal evolution of the $B_{\mathrm{rms}}$ and diffusivity $\eta$ as a function of the radius for the $0.6\ M_{\odot}$ model assuming a dipole geometry and taking $\mathrm{f_{RM}}=10$ for turbulent diffusivity. This figure is similar to Figure 3 of Castro-Tapia et al. (2024b). However, in the current case, we note how the magnetic field has already diffused to the surface of the WD when less than $5\%$ of the crystal core has formed as a direct consequence of the initial condition of the magnetic boundary computed by Camisassa et al. (2024) and letting the magnetic field evolve from the formation of the WD.
In order to later compare with observations, we computed the surface value of $B_{\mathrm{rms}}$ as a function of time from the magnetic field evolution of the models with different masses. In the left panel of Figure 4, we show $B_{\mathrm{rms,surf}}$ normalized to the initial value $B_{0}$, as a function of the cooling time when considering only ohmic diffusion for the different masses and d and d+q field geometries. We define the reduction factor $f_{B}(t_{\mathrm{cool}})=(B_{\mathrm{rms,sur}}/B_{0})(t_{\mathrm{cool}})$, which indicates the expected value of the magnetic field at the WD surface, normalized by the initial field, as a function of the WD cooling time. In the right panel, we show the same but including turbulent diffusion with $\mathrm{f_{RM}}=10$. Regardless of the mass, all the cases show that $B_{\mathrm{rms,surf}}$ increases as it is transported from the interior to the surface, reaching a maximum intensity, to then decay, as it is diffused out of the WD, with a much smaller rate of change than the first increase. This slow decay is a direct consequence of the crystallization process that increases the conductivities as the solid phase grows.
@@ III.2 Magnetic field evolution for different masses
IV Comparison with observations
Figure 5: Color map indicating the logarithmic prediction of the magnetic field at the WD surface as a function of the WD mass and cooling time. For this, we assumed dipole geometry, a $B_{0}$ scaled to equipartition from the main-sequence dynamo, and considered magnetic flux conservation. For comparison, the non-MWDs and MWDs from the local 20-40pc sample of Bagnulo & Landstreet (2022) are shown with the same color scale as the background for their measured field. The crystallization onset and the moment when the magnetic field models reach their maximum as a function of the WD mass are plotted using black and cyan lines, respectively (the dashed part of the cyan line takes the same age as the maximum of the $1\ M_{\odot}$ model for the more massive WDs).
@@ IV Comparison with observations
Figure 6: Left panel: Magnetic fields at the surface of white dwarf stars taken from the Montreal White Dwarf Database (red crosses), and their corresponding magnetic fields expected at the edge of the magnetic boundary during the main sequence stage, assuming different considerations for the field geometry and diffusivities (circles). The solid lines indicate the magnetic fields expected at this magnetic boundary during the main sequence, taken from the MHD simulations of main sequence A stars of Brun et al. (2005) (green lines) and Hidalgo et al. (2024) (black line) and from our estimations assuming equipartition between the kinetic and magnetic energy (purple line). The blue shaded area indicates the typical magnetic field strengths measured at the surface of magnetic main sequence Ap and Bp stars. Right panel: Same as left panel but for the sample of magnetic white dwarfs within 20pc from our Sun and magnetic white dwarfs younger than 0.6 Gyrs within 40pc from our Sun (Bagnulo & Landstreet, 2022).
@@
Figure 7: Predictions of the surface magnetic field strengths (rms) as a function of cooling age for our models compared to observed magnetic fields in a 100 pc sample of the MWDD combined with the MWDs of Bagnulo & Landstreet (2022). The initial magnetic field prediction $B_{0}$ comes from equipartition in the main-sequence and rescaling assuming conservation of the magnetic flux (Eq. 6). We have separated into different panels the masses of the models in the range $0.6$-$1\ M_{\odot}$, and we compare in each panel with MWDs with masses within a range of $\pm 0.05\ M_{\odot}$ for the corresponding model mass. The solid lines show the d geometry models, while the dashed lines are the d+q models. For each geometry, the split in the lines at some point of the evolution shows the effect of turbulent diffusion with $\mathrm{f_{R_{m}}}=10$, being the cases of more rapid decay the ones that include this factor. In the bottom-right panel, we summarize the range of the magnetic field strengths predicted and observed for different masses; for the observations used in the previous panels we show a logarithmic average of the magnetic fields in the bin, whereas for the model range, the upper limit is the maximum strength predicted, and the arrow is at the minimum $B_{\mathrm{rms,surf}}$ for the age of the youngest WD in the bin. A linear fit in the logarithmic space of mass and B (magenta line) for all the data points shown (in black) is included. The shaded (magenta) area shows the $\pm 5\sigma$ range of the fit parameters fitted (see text).
To compare our magnetic field evolution models with observed surface magnetic fields in WDs, we take the local 20-40 pc sample of WDs from Bagnulo & Landstreet (2022) and the MWDs available in the Montreal White Dwarf Database (MWDD, Dufour et al., 2017). In Figure 5, we show the mass and cooling age of the MWDs from the 20-40 pc sample. As a background color map, we show the logarithm of the $B_{\mathrm{rms,surf}}$ value for each mass and cooling age predicted by our dipole evolution models, where we assumed a $B_{0}$ for each mass scaled according to magnetic flux conservation as $B_{0}=B_{\mathrm{MS}}(R_{\mathrm{WD,b}}/R_{\mathrm{MS,b}})^{2}$, where $R_{\mathrm{MS,b}}$ and $R_{\mathrm{WD,b}}$ are the radius at the magnetic boundary of the main-sequence progenitor and the WD, respectively, and taking the equipartition prediction for $B_{\mathrm{MS}}$ [^1] As we only computed models for the magnetic field evolution of up to $1\ M_{\odot}$ WDs, for the colormap background of $1\ M_{\odot}$ we used the same reduction factor predicted for the $1\ M_{\odot}$ case and extrapolated $B_{0}$ using a linear fit in the logarithmic space with the mass.. The same color scaling is used to show the value of the observed surface magnetic field of the MWDs in the sample. This comparison confirms the results of Camisassa et al. (2024): the late appearance of magnetic fields in WDs is not necessarily related to the crystallization process. For any fixed mass in Figure 5, we see that the MWDs tend to accumulate around the point where the maximum of the model prediction is reached. We also see that most of the magnetic WDs are located in the regions where we expect the surface magnetic field to be larger. While the majority of MWDs appear near or after the onset of crystallization, those that appear before can only be explained by the appearance of a magnetic field formed before the WD phase, and may not be related to crystallization at all.
While the models shown in Figure 5 predict that for increasing mass, more MWDs must appear over a wide range of cooling age, we acknowledge that our models consider an ideal case, where the magnetic field generated during the main-sequence, with perfect equipartition assumption, reaches the WD phase without energy loss. Despite these caveats, we see that at about >0.1>0.1 Gyr, for the $\gtrsim 1\ M_{\odot}$ models, the field predictions are of the same order of magnitude as the maximum magnetic field reached. Something similar happens for the less massive MWDs, but at $\gtrsim 1$ Gyr. Moreover, following the color scaling, it seems that the observations are almost all at the same or below the value predicted by the background map, confirming that the models can be interpreted as the best-case scenario for the magnetic field formation and evolution.
@@
In order to predict the value of the magnetic field in the main-sequence phase $B_{\mathrm{MS}}$ necessary to reproduce the observed surface fields in MWDs we consider the reduction factors $f_{B}(t_{\mathrm{cool}})$ from the different evolution models and the assumption of magnetic flux conservation given by $R_{\mathrm{MS,b}}$ and $R_{\mathrm{WD,b}}$. Thus, for a given observed field $B_{\mathrm{rms,surf}}$ in the WD phase, its corresponding magnetic field strength generated during the main-sequence can be computed as follows
@@
where $f_{B}(t_{\mathrm{cool}})$ is the prediction of our model at a given cooling time $t_{\mathrm{cool}}$ because in this case we do not need to know $B_{0}$ a priori.
@@
In Figure 6 we show the observed magnetic field in MWDs (red crosses) for different masses and their corresponding predictions of the main-sequence field $B_{\mathrm{MS}}$ (empty circles) necessary to reproduce the observations, obtained using equation (6) for our models with d and d+q geometries. We also included the prediction from our models with turbulent diffusion. In the left panel of the figure, we show the MWDs in the MWDD, while the right panel shows the MWDs in the 20-40 pc sample of Bagnulo & Landstreet (2022). To compare our $B_{\mathrm{MS}}$ predictions, we added the magnetic field strength prediction from a linear fit to the equipartition value at the magnetic boundary of main-sequence progenitors for different WDs, assuming an initial-to-final-mass relation from (Catalán et al., 2008). We also compare the predicted magnetic field strengths from core convection in the main-sequence obtained from MHD simulations of Brun et al. (2005) and Hidalgo et al. (2024), for Main Sequence masses of 2 and 2.2M, respectively.
@@
We note that the necessary $B_{\mathrm{MS}}$ predictions to reproduce the observations are mostly in the range $\sim 10^{3}$-$10^{5}$ G, which is very similar to the equipartition and MHD predictions of $\sim 10^{4}$-$10^{5}$ G. The equipartition prediction appears almost as an upper limit for the main-sequence convective core dynamo field necessary to reproduce the observed MWDs with $\gtrsim 0.75\ M_{\odot}$, except for a very few whose $B_{\mathrm{MS}}$ predictions are larger than the equipartition by a factor of $\lesssim 5$. For MWDs with $\lesssim 0.75\ M_{\odot}$, the $B_{\mathrm{MS}}$ predictions appear more scattered from the equipartition value, some of them even being the highest $B_{\mathrm{MS}}$ for all the masses. This occurs because of the small magnetic boundaries and larger radius for the less massive WDs, which lead to a maximum factor $f_{B}\lesssim 0.03$. Therefore, to reproduce MWDs with about the same few MG strength, the $B_{\mathrm{MS}}$ must be larger for the less massive stars. However, for all the masses, about half of the magnetic fields observed in WDs could have originated from $B_{\mathrm{MS}}$ weaker than $10^{4}$ G, below the equipartition and MHD simulations prediction. Some MWDs with $10^{5}$-$10^{6}$ G could have even been produced with $B_{\mathrm{MS}}$ of just a hundred G. It is important to remark that the initial-to-final-mass relation and, consequently, the value of $R_{\mathrm{MS,b}}/R_{\mathrm{WD,b}}$ are subject to large uncertainties, particularly for the more massive WDs. Therefore, the value of $B_{\mathrm{MS}}$ could be much weaker or stronger than predicted here.